The B/C ratio for a flood control project along the Swanee River was calculated to be 1.5. If the benefits were $450,000 per year and the maintenance costs were $221,000 per year, determine the initial cost of the project at an interest rate of 7% per year and a 50-year life.

The initial cost of the project is
Answer #1

Let the initial cost be X

B/C ratio = PW of benefits/PW of costs

PW of benefits= 450000(P/A,7%,50) = 450000(13.80075) = 6210336

PW of disbenefits = X+221000(P/A,7%,50) = X+3049965

So,

1.5 = 6210336/X+3049965

1.5X+4574947.5 = 6210336

X= 1635388.5/1.5

X= 1090259